OpenStudy (anonymous):

Hi I need some help with (Modeling with Functions) The total daily cost (in dollars) of manufacturing scented candles is given by the function C(b) = 250 + 3b, where b is the number of candles manufactured. Which function represents the average manufacturing cost, A, in terms of b?

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

could you help me @mathmath333

OpenStudy (mathmath333):

where are the options

OpenStudy (anonymous):

1. A(b) = b/250+3b 2. A(b) = 250+3b/b 3. A(b) = (250+3b) x (b) 4. A(b) = 250+4b

OpenStudy (anonymous):

i think it is either 3 or 4

OpenStudy (mathmath333):

is options be this \(\large \color{black}{\begin{align} A(b)=250+\dfrac{3b}{b} \hspace{1.5em}\\~\\ \end{align}}\) or this \(\large \color{black}{\begin{align} A(b)=\dfrac{ 250+3b}{b} \hspace{1.5em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

where there is a / im trying to say that everything befor that is above it

OpenStudy (mathmath333):

this is option B ? \(\large \color{red}{\begin{align} A(b)=\dfrac{250+3b}{b} \hspace{1.5em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

the average can be calculated when we divided the whole by number of times it occurs so it should be B

OpenStudy (anonymous):

OH! ok now i understand, thanks! that helped alot! !